OpenStudy (anonymous):

evaluate limit: lim ((x+h)^3 -3)/h h->0

OpenStudy (anonymous):

is that x+h or 3+h?

OpenStudy (anonymous):

it is the quantity (x+h) to the third power minus 3. all over h

OpenStudy (anonymous):

\[(x+h)^3=x^3+3x^2h+3xh^2+h^3\]

OpenStudy (anonymous):

there is no limit because you get \[\frac{x^3+3x^2h+3xh^2+h^3-3}{h}\] nothing factors, nothing cancels. which is why i am wondering if there is a typo because it looks like you are trying to find a derivative, but something is amiss

OpenStudy (anonymous):

are you sure it isn't \[\lim_{h\rightarrow 0}\frac{(x+h)^3-x^3}{h}\]

OpenStudy (anonymous):

yes thats it. sorry if i was confusing

OpenStudy (anonymous):

how did i guess?

OpenStudy (anonymous):

\[\frac{(x+h)^3-x^3}{h}=\frac{x^3+3x^2h+3xh^2+h^2-x^3}{h}=\frac{3x^2h+3xh^2+h^3}{h}\] \[=\frac{h(3x^2+3xh+h^2}{h}=3x^2+3xh+h^2\]

OpenStudy (anonymous):

take the limit as h goes to zero get \[3x^2\] and in a week you will be able to do this instantly in your head

OpenStudy (anonymous):

wow thank you, could you help me on the other problem i posted??

OpenStudy (anonymous):

i will find it